what is the equation of the line that passes through the points (-2,2) and (0,5)?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

what is the equation of the line that passes through the points (-2,2) and (0,5)?
----------------
slope = (5-2)/(0--2) = 3/2
--------------
intercept: (0,5)
---------------
EQUATION:
y = (3/2)x + 5
====================
Cheers,
Stan H.

Answer by Alan3354(69443)

what is the equation of the line that passes through the points (-2,2) and (0,5)?
--------
|x y 1|
|-2 2 1| = 0
|0 5 1|
x*(2-5) - y*(-2) + -10 = 0
-3x + 2y -10 = 0
3x - 2y = -10